# Prox-DBRO-VR: A Unified Analysis on Decentralized Byzantine-Resilient Composite Stochastic Optimization with Variance Reduction and Non-Asymptotic Convergence Rates

### v6

@@ VI Conclusions
We define ${U_{k}}:=\left\|{{x_{k}}-{x^{*}}}\right\|_{2}^{2}+{q_{\min}}\gamma{\alpha_{k}}t_{k}/L$, which is non-negative, since $t_{k}$ is non-negative. We further set $0<{\alpha_{k}}\leq 4\gamma/\left({{\kappa_{q}}\left({64{L^{2}}+{q_{\min}}{\gamma^{2}}}\right)}\right)$ and take the full expectation on the both sides of (72) to obtain
${\mathbb{E}}\left[{{U_{k+1}}}\right]\leq$ ${\mathbb{E}}\left[{\left\|{{x_{k+1}}-{x^{*}}}\right\|_{2}^{2}}\right]+\frac{{{q_{\min}}\gamma{\alpha_{k}}}}{L}{\mathbb{E}}\left[{t_{k+1}}\right]$ (73)
$\leq$ $\left({1-\frac{\gamma}{4}{\alpha_{k}}}\right){\mathbb{E}}\left[{U_{k}}\right]+P_{1}^{d}\alpha_{k}^{2}+{P_{2}}{\alpha_{k}},$
where the first inequality is due to the fact that the step-size $\alpha_{k}$ is decaying. Via summarizing all the upper bounds on the decaying step-size, we attain a feasible selection range as follows:
@@
$\mathbb{E}\left[{{U_{k}}}\right]\leq\Xi/\left({k+\xi}\right)+\tilde{E},\forall k\geq 0,$ (75)
@@
${U_{1}}\leq\left({1-\frac{\gamma}{4}{\alpha_{0}}}\right){U_{0}}+\alpha_{0}^{2}{P_{1}^{d}}+{\alpha_{0}}{P_{2}}.$ (76)
Therefore, for a bounded and positive constant $\tilde{E}$, if $\Xi\geq\left({\xi-\gamma\theta/4}\right){U_{0}}+{{\theta^{2}}{P_{1}^{d}}/\xi}+\theta{P_{2}}-\xi\tilde{E}$, we have
$U_{1}\leq\left({1-\frac{\gamma}{4}{\alpha_{0}}}\right){U_{0}}+\alpha_{0}^{2}{P_{1}^{d}}+{\alpha_{0}}{P_{2}}\leq\frac{\Xi}{\xi}+\tilde{E},$ (77)
with ${\alpha_{0}}=\theta/\xi$. We assume that for $k=k^{\prime}$, $k^{\prime}\geq 1$, it satisfies that
$\mathbb{E}\left[{{U_{k^{\prime}+1}}}\right]\leq$ $\left({1-\frac{\gamma}{4}{\alpha_{k^{\prime}}}}\right)\mathbb{E}\left[{{U_{k^{\prime}}}}\right]+\alpha_{k^{\prime}}^{2}{P_{1}}+{\alpha_{k^{\prime}}}{P_{2}}$ (78)
$\leq$ $\frac{\Xi}{{k^{\prime}+\xi}}+\tilde{E}.$
Then, we will prove that for $k=k^{\prime}+1$,
$\mathbb{E}\left[{{U_{k^{\prime}+2}}}\right]\leq\frac{\Xi}{{k^{\prime}+\xi+1}}+\tilde{E},$ (79)
holds true. Define $\tilde{\gamma}:=\gamma/4$ and set $\tilde{E}\geq{P_{2}}/\tilde{\gamma}$ and $\Xi\geq{\theta^{2}}{P_{1}}/\left({\tilde{\gamma}\theta-1}\right)$ with $\theta>1/\tilde{\gamma}$. We have
$\mathbb{E}\left[{{U_{k^{\prime}+2}}}\right]$ (80)
$\leq$ $\left({1-\tilde{\gamma}{\alpha_{k^{\prime}+1}}}\right)\mathbb{E}\left[{{U_{k^{\prime}+1}}}\right]+\alpha_{k^{\prime}+1}^{2}{P_{1}}+{\alpha_{k^{\prime}+1}}{P_{2}}$
$\leq$ $\left({1-\frac{{\tilde{\gamma}\theta}}{{k^{\prime}+\xi+1}}}\right)\left({\frac{\Xi}{{k^{\prime}+\xi}}+\tilde{E}}\right)+\frac{{{\theta^{2}}}}{{{{\left({k^{\prime}+\xi+1}\right)}^{2}}}}{P_{1}}$
$+\frac{\theta}{{k^{\prime}+\xi+1}}{P_{2}}$
$\leq$ $\left({1-\frac{{\tilde{\gamma}\theta}}{{k^{\prime}+\xi+1}}}\right)\frac{\Xi}{{k^{\prime}+\xi}}+\tilde{E}+\frac{{{\theta^{2}}}}{{{{\left({k^{\prime}+\xi+1}\right)}^{2}}}}{P_{1}}$
$\leq$ $\left({1-\frac{{\tilde{\gamma}\theta}}{{k^{\prime}+\xi+1}}}\right)\frac{\Xi}{{k^{\prime}+\xi}}+\tilde{E}+\frac{{\Xi\left({\tilde{\gamma}\theta-1}\right)}}{{{{\left({k^{\prime}+\xi+1}\right)}^{2}}}}$
$\leq$ $\left({1-\frac{{\tilde{\gamma}\theta}}{{k^{\prime}+\xi+1}}}\right)\frac{\Xi}{{k^{\prime}+\xi}}+\frac{{\Xi\left({\tilde{\gamma}\theta-1}\right)}}{{\left({k^{\prime}+\xi+1}\right)\left({k^{\prime}+\xi}\right)}}+\tilde{E}$
$=$ $\frac{\Xi}{{k^{\prime}+\xi+1}}+\tilde{E},$
which means the relation (75) holds true. Via replacing $\tilde{E}$ with its lower bound $E={P_{2}}/\tilde{\gamma}$, it is straightforward to verify

### v12

@@ VI Conclusion
We define ${\tilde{U}_{k}}:=\left\|{{x_{k}}-{x^{*}}}\right\|_{2}^{2}+{q_{\min}}\gamma{\alpha_{k}}t_{k}/L$, which is non-negative, since $t_{k}$ is non-negative. We further set $0<{\alpha_{k}}\leq 4\gamma/\left({{\kappa_{q}}\left({64{L^{2}}+{q_{\min}}{\gamma^{2}}}\right)}\right)$ and take the total expectation on the both sides of (73) to obtain
${\mathbb{E}}\left[{{\tilde{U}_{k+1}}}\right]\leq$ ${\mathbb{E}}\left[{\left\|{{x_{k+1}}-{x^{*}}}\right\|_{2}^{2}}\right]+\frac{{{q_{\min}}\gamma{\alpha_{k}}}}{L}{\mathbb{E}}\left[{t_{k+1}}\right]$ (74)
$\leq$ $\left({1-\frac{\gamma}{4}{\alpha_{k}}}\right){\mathbb{E}}\left[{\tilde{U}_{k}}\right]+P_{1}^{d}\alpha_{k}^{2}+{P_{2}}{\alpha_{k}},$
where the first inequality is due to the fact that the step-size $\alpha_{k}$ is decaying. By summarizing all the required upper bounds on the step-size, it suffices to consider a feasible range as follows:
@@
$\mathbb{E}\left[{{\tilde{U}_{k}}}\right]\leq\Xi/\left({k+\xi}\right)+\tilde{E},\forall k\geq 0,$ (76)
@@
${\tilde{U}_{1}}\leq\left({1-\frac{\gamma}{4}{\alpha_{0}}}\right){\tilde{U}_{0}}+\alpha_{0}^{2}{P_{1}^{d}}+{\alpha_{0}}{P_{2}}.$ (77)
Therefore, for a bounded and positive constant $\tilde{E}$, if $\Xi\geq\left({\xi-\gamma\theta/4}\right){\tilde{U}_{0}}+{{\theta^{2}}{P_{1}^{d}}/\xi}+\theta{P_{2}}-\xi\tilde{E}$, we have
$\tilde{U}_{1}\leq\left({1-\frac{\gamma}{4}{\alpha_{0}}}\right){\tilde{U}_{0}}+\alpha_{0}^{2}{P_{1}^{d}}+{\alpha_{0}}{P_{2}}\leq\frac{\Xi}{\xi}+\tilde{E},$ (78)
with ${\alpha_{0}}=\theta/\xi$. We assume that for $k=K$, $K\geq 1$, it satisfies that
$\mathbb{E}\left[{{\tilde{U}_{K+1}}}\right]\leq$ $\left({1-\frac{\gamma}{4}{\alpha_{K}}}\right)\mathbb{E}\left[{{\tilde{U}_{K}}}\right]+\alpha_{K}^{2}{P_{1}}+{\alpha_{K}}{P_{2}}$ (79)
$\leq$ $\frac{\Xi}{{K+\xi}}+\tilde{E}.$
Then, we will prove that for $k=K+1$,
$\mathbb{E}\left[{{\tilde{U}_{K+2}}}\right]\leq\frac{\Xi}{{K+\xi+1}}+\tilde{E},$ (80)
holds true. We define $\tilde{\gamma}:=\gamma/4$ and set $\tilde{E}\geq{P_{2}}/\tilde{\gamma}$ and $\Xi\geq{\theta^{2}}{P_{1}}/\left({\tilde{\gamma}\theta-1}\right)$ with $\theta>1/\tilde{\gamma}$. We have
$\mathbb{E}\left[{{\tilde{U}_{K+2}}}\right]$ (81)
$\leq$ $\left({1-\tilde{\gamma}{\alpha_{K+1}}}\right)\mathbb{E}\left[{{\tilde{U}_{K+1}}}\right]+\alpha_{K+1}^{2}{P_{1}}+{\alpha_{K+1}}{P_{2}}$
$\leq$ $\left({1-\frac{{\tilde{\gamma}\theta}}{{K+\xi+1}}}\right)\left({\frac{\Xi}{{K+\xi}}+\tilde{E}}\right)+\frac{{{\theta^{2}}}}{{{{\left({K+\xi+1}\right)}^{2}}}}{P_{1}}$
$+\frac{\theta}{{K+\xi+1}}{P_{2}}$
$\leq$ $\left({1-\frac{{\tilde{\gamma}\theta}}{{K+\xi+1}}}\right)\frac{\Xi}{{K+\xi}}+\tilde{E}+\frac{{{\theta^{2}}}}{{{{\left({K+\xi+1}\right)}^{2}}}}{P_{1}}$
$\leq$ $\left({1-\frac{{\tilde{\gamma}\theta}}{{K+\xi+1}}}\right)\frac{\Xi}{{K+\xi}}+\tilde{E}+\frac{{\Xi\left({\tilde{\gamma}\theta-1}\right)}}{{{{\left({K+\xi+1}\right)}^{2}}}}$
$\leq$ $\left({1-\frac{{\tilde{\gamma}\theta}}{{K+\xi+1}}}\right)\frac{\Xi}{{K+\xi}}+\frac{{\Xi\left({\tilde{\gamma}\theta-1}\right)}}{{\left({K+\xi+1}\right)\left({K+\xi}\right)}}+\tilde{E}$
$=$ $\frac{\Xi}{{K+\xi+1}}+\tilde{E},$
which means the relation (76) holds true. Via replacing $\tilde{E}$ with its upper bound $E={P_{2}}/\tilde{\gamma}$, it is straightforward to verify